# Modality-Agnostic Structural Image Representation Learning for Deformable Multi-Modality Medical Image Registration

### v1

@@ 1 Introduction
  • we introduce a Deep Neighbour Self-similarity (DNS) module, which can capture long-range and complex structural information from medical images, addressing the ambiguity in classical feature descriptors and similarity metrics.
@@ 2 Related Work
Prominent examples of statistical-based methods use information theory and optimize joint voxel statistics, such as minimizing the MI or normalized mutual information (NMI) as similarity measures for multi-modal registration [3, 4, 52]. These approaches aim to estimate a solution that minimizes the entropy of the joint histogram between image pairs for rigid registration. However, as evidenced in [45], MI-based similarity measure is restricted to measuring the statistical co-occurrence of image intensities, which are not extendable to the co-occurrence of complex patterns, such as subtle structural information of soft tissues and vascular structures in medical images, and are sensitive to locally varying noise distribution, which is not ideal for deformable multi-modal registration.
[…]
@@ 3 Method
Figure 2: Overview of the proposed Modality-Agnostic Deep Structural Representation Network (MASR-Net) and anatomy-aware contrastive learning paradigm. For brevity, we visualize the 3D feature maps in a 2D aspect. Only the negative pairs of the first feature vector are shown.
@@ 3.2 Modality-Agnostic Deep Structural Representation Network
We first leverage a feed-forward 3D convolutional neural network (CNN) to extract the image feature $\bm{h}$ for the input image $I$. The proposed CNN network is built with a 4-level encoder-decoder structure with skip connection [47], which is composed of 3D convolution layers, LeakyReLU activations [31] and uses BlurPool [61] and trilinear interpolation for downsampling and upsampling, respectively, resulting in a maximum striding factor of 8. The network takes the input image $I\in\mathbb{R}^{C\times H\times W\times D}$ and outputs an image feature map $\bm{h}\in\mathbb{R}^{C_{h}\times H\times W\times D}$. The architectural details of the network are shown in the supplementary material.
@@
$\begin{array}[]{l}\!\mathbf{S}(\bm{h},x,y)\!=\!\exp\Big{(}\!-\!\sum_{y^{\prime}\in\mathcal{N}(x)}\frac{(\bm{h}(y)-\bm{h}(y^{\prime}))^{2}}{\sigma^{2}}\Big{)},y^{\prime}\neq y\end{array}$ (1)
where $y,y^{\prime}\in\mathcal{N}(x)$ defines the neighbour location of $x$ and $|\mathcal{N}(x)|$ represents the number of neighbour voxels in $\mathcal{N}(x)$. The denominator $\sigma^{2}$ is a noise estimator, defined as the mean of all patch distances, i.e., $\frac{1}{|\mathcal{N}(x)|}\sum_{y^{\prime}\in\mathcal{N}(x)}(\bm{h}(y)-\bm{h}(y^{\prime}))^{2}$. We follow [20] to further restrict the pairwise distance calculations within the six-neighbourhood (6-NH) with a Euclidean distance of $\sqrt{2}$ between them, reducing the computation complexity of DNS, i.e., reduce the number of unique pair-wise distances from $15$ to $12$.
@@ 3.3 Anatomy-aware Contrastive Learning
$\begin{array}[]{l}B(t)=\sum_{i=0}^{n}P_{i}\cdot b_{i,n}(t)\end{array}$ (2)
@@
$I^{\prime}(x)=\begin{cases}B(I(x)),&\text{if }p>\delta\\ B(1-I(x)),&\text{if }p\leq\delta\end{cases}\quad x\in\Omega$ (3)
@@ 3.4 Multimodal Similarity Using DNS
$\begin{array}[]{l}\!\phi^{*}\!=\!\operatorname*{arg\,min}_{\phi_{\theta}}\mathcal{L}_{sim}(\psi_{\sigma}(\mathbf{D}_{F}),\psi_{\sigma}(\mathbf{D}_{M}(\phi_{\theta})))\!+\!\lambda\mathcal{L}_{reg}(\phi_{\theta}),\end{array}$ (5)
where $\phi^{*}$ denotes the optimal displacement field, the $\psi_{\sigma}$ is a Gaussian smoothing function with variance $\sigma$, $\mathcal{L}_{sim}(\mathbf{D}_{F},\mathbf{D}_{M}(\phi_{\theta}))$ denotes the negated cosine similarity function that quantifies the dissimilarity between the $\mathbf{D}_{F}$ and the $\mathbf{D}_{M}(\phi_{\theta})$ spatially transformed by $\phi$, and $\mathcal{L}_{reg}(\phi)=||\nabla\phi||^{2}_{2}$ represents the smoothness regularization function that penalizes implausible solution. Our proposed DNS is differentiable and can be adapted to iterative instance optimization methods [40, 51] and learning-based registration frameworks [6, 37, 35, 36]. To exemplify the flexibility of DSIR, we parameterize the registration function with one learning-based method, LapIRN [37], and one 3-level non-rigid instance optimization method (IO) [38]. The implementation details of LapIRN and IO are shown in the supplementary material.
@@ 4.1 Data and Pre-processing
We collected a 3D multi-phase contrast-enhanced liver CT dataset from hospitals in China to evaluate the inter-patient registration of liver CTs. This dataset consists of 1966 patients with tumors, and each patient contains scans in three phases, i.e., pre-contrast, arterial and venous phases, acquired from the CT scanner at different time points after the contrast agent injection. There are large non-linear deformations and heterogeneous image intensities across multi-phase CT due to gravity, muscle contractions, and contrast agent injection. We divided the dataset into 1886/10/70 cases for training, validation, and test sets. The liver and tumour masks of the validation and test sets are manually annotated and checked by a senior radiologist.
@@
We also evaluated our method on the Abdomen MR-CT task of the Learn2Reg challenge 2021 [22]. The dataset comprises 8 sets of paired MR and CT scans and 90 unpaired CT/MR scans, both depicting the abdominal region of a single patient and presenting notable deformations. Since there are only limited paired MR and CT scans, we only evaluate the iterative optimization-based registration methods in this task. We use all 8 paired MR and CT scans as the test set. Each scan has four manually labelled anatomical structures: liver, spleen, right kidney, and left kidney. For MASR-Net, we split the 90 unpaired CT/MR scans into 85 and 5 scans for training and validation sets.
@@
[…]
@@ 4.3 Evaluation metrics and Baselines
We compare our method with three state-of-the-art conventional registration methods (ANTs [4], NiftyReg [52] and DEEDs [18]) and two learning-based registration approaches (Voxelmorph [6] and LapIRN [41]). Specifically, we use the ANTs registration implementation in the publicly available ANTs software package [3]. Both ANTs and NiftyReg methods use a four-level multi-resolution optimization strategy with adaptive gradient descent optimization and MI as the similarity measure. For Voxelmorph and LapIRN, we use their official implementations and the best hyperparameter reported in their paper. All learning-based methods are trained with two similarity metrics (NMI and MIND) by default.
@@ 4.4 Results
Flexibility and versatility. Compared to the results of LapIRN with different metrics, LapIRN with our proposed DNS achieves the best overall registration performance in tumor and organ alignment, showing +2.22% and +0.75 DSC gains over the one with MIND in tumor and organ alignment, respectively. More importantly, when comparing the results of IO with DNS and MIND, there are significant improvements in registration accuracy of both tumor and organ (+4.16 and + 1.72 gains in DSC, respectively). The result proves that our method is more robust and flexible with diverse registration frameworks than MIND. When using DNS with a simple iterative gradient optimization, the DSIR of DNS is more discriminative than MIND and can lead to less ambiguous matching during the iterative optimization. Besides, although NMI shows comparable DSC in multiphase CT registration, the qualitative results in Fig. 3 demonstrate that methods using MI/NMI fail to recover the large non-linear misalignment in MR-CT registration.
@@
Robust to large deformation. The initial DSC of this task is relatively low (37.32), indicating there are possible large spatial misalignments between image pairs. All conventional methods (ANTs and NiftyReg) with MI fail spectacularly in this task, suggesting that MI cannot register multimodal images with large deformation. In contrast to MI, structural image representation-based methods (DEEDs and DNS) achieve decent results in this task (+83.53 and +85.61 DSC, respectively). It is worth noting that our DNS with a solely simple gradient descent-based optimization strategy (IO) outperforms DEEDs, which employ a more sophisticated convex optimization, in terms of registration accuracy and runtime in this task, further suggesting the high discriminability of our DSIR.
@@ 4.4.1 Ablation Studies
Effect of Deep Neighbourhood Self-similarity and contrastive learning. Tab. 3 demonstrates performance metrics for different registration models on the multiphase CT registration task (Pre-contrast $\leftarrow$ Venous & Arterial). Although a recent study [13] reported that random features from CNN could be expressive, the DSIR of random initialize CNN has lower DSC than that of initial, suggesting that random initialization alone is insufficient for multi-modal registration. Moreover, there is a significant performance gain when the DNS module is added to the randomly initialized CNN (+11.12 DSC), suggesting the effectiveness of DNS in extracting deep structural information. Comparing the feature similarity analysis of random init. (MASR-Net without training) and DNS in Fig. 4, one can observe that the proposed anatomy-aware contrastive learning effectively improves the discriminability of the DSIR of DNS, resulting in high feature similarity centralized in the true anatomical correspondence and relatively low feature similarity at the other anatomical structures.
@@ 5 Conclusion
This paper proposes a modality-agnostic structural image representation learning method dedicated to multi-modality medical image registration. Our method leverages deep neighbourhood self-similarity to learn highly discriminative, contrast invariance structural representations for multimodal images. Anatomy-aware contrastive learning is introduced to further enhance the expressiveness and discriminability of the learned structural image representation, reducing the ambiguity in matching the true anatomical correspondence between multimodal images. Comprehensive experiments demonstrate that our method effectively reduces the multi-modality registration problem to a monomodal one, achieving state-of-the-art registration accuracy compared with conventional local structural representations without the need for a dedicated optimizer.

### v2

@@ 1 Introduction
  • we introduce the Deep Neighbour Self-similarity (DNS), which can capture long-range and complex structural information from medical images, addressing the ambiguity in classical feature descriptors and similarity metrics.
@@ 2 Related Work
Prominent examples of statistical-based methods use information theory and optimize joint voxel statistics, such as minimizing the MI or normalized mutual information (NMI) as similarity measures for multi-modal registration [3, 4, 52]. These approaches aim to estimate a solution that minimizes the entropy of the joint histogram between image pairs for rigid registration. However, as evidenced in [45], MI-based similarity measures are restricted to measuring the statistical co-occurrence of image intensities, which are not extendable to the co-occurrence of complex patterns, such as subtle structural information of soft tissues and vascular structures in medical images, and are sensitive to locally varying noise distribution, which is not ideal for deformable multi-modal registration.
To circumvent the limitation of statistical-based methods, another common strategy is to project multimodal images into common intermediate structural representations or measure the misalignment through the modality-invariant similarity metrics [27, 29, 12]. Prominent examples include MIND  [17, 18, 20], attribute vectors [49] and Linear Correlation of Linear Combination (LC2) [56] that are designed to capture a dense structural image representation of multi-modal images that are independent of the underlying image acquisition. However, these local structural representations are often not sufficiently discriminative and expressive to drive a non-rigid registration with many degrees of freedom, not differentiable or expensive to compute. As such, these local structural representations are often used in conjunction with robust optimization methods, e.g. discrete optimization [19] and Bound Optimization by Quadratic Approximation [46], or requiring robust rigid initialization.
@@ 3 Method
Figure 2: Overview of the Modality-Agnostic Deep Structural Representation Network (MASR-Net) and anatomy-aware contrastive learning paradigm. For brevity, we visualize the 3D feature maps in a 2D aspect. Only negative pairs of the first feature vector are shown.
@@ 3.2 Modality-Agnostic Deep Structural Representation Network
We first leverage a feed-forward 3D convolutional neural network (CNN) to extract the image feature $\bm{h}$ for the input image $I$. The proposed CNN network is built with a 4-level encoder-decoder structure with skip connection [47], which is composed of 3D convolution layers, LeakyReLU activations [31] and uses BlurPool [61] and trilinear interpolation for downsampling and upsampling, respectively, resulting in a maximum striding factor of 8. The network takes the input image $I\in\mathbb{R}^{C\times H\times W\times D}$ and outputs an image feature map $\bm{h}\in\mathbb{R}^{C_{h}\times H\times W\times D}$. The details of the network are shown in supplementary material.
@@
$\begin{array}[]{l}\!\mathbf{S}(\bm{h},x,y)\!=\!\exp\Big{(}\!-\!\sum_{y^{\prime}\in\mathcal{N}(x)}\frac{(\bm{h}(y)-\bm{h}(y^{\prime}))^{2}}{\sigma^{2}}\Big{)},y^{\prime}\neq y,\end{array}$ (1)
where $y,y^{\prime}\in\mathcal{N}(x)$ defines the neighbour location of $x$. The denominator $\sigma^{2}$ is a noise estimator, defined as the mean of all patch distances, i.e., $\frac{1}{|\mathcal{N}(x)|}\sum_{y^{\prime}\in\mathcal{N}(x)}(\bm{h}(y)-\bm{h}(y^{\prime}))^{2}$, where $|\mathcal{N}(x)|$ represents the number of neighbour voxels in $\mathcal{N}(x)$. We follow [20] to further restrict the pairwise distance calculations within the six-neighbourhood (6-NH) with a Euclidean distance of $\sqrt{2}$ between them, reducing the computation complexity of DNS, i.e., reduce the number of unique pair-wise distances from $15$ to $12$.
@@ 3.3 Anatomy-aware Contrastive Learning
$\begin{array}[]{l}B(t)=\sum_{i=0}^{n}P_{i}\cdot b_{i,n}(t),\end{array}$ (2)
@@
$I^{\prime}(x)=\begin{cases}B(I(x)),&\text{if }p>\delta\\ B(1-I(x)),&\text{if }p\leq\delta\end{cases}\quad x\in\Omega,$ (3)
@@ 3.4 Multimodal Similarity Using DNS
$\begin{array}[]{l}\!\phi^{*}\!=\!\operatorname*{arg\,min}_{\theta}\mathcal{L}_{sim}(\psi_{\sigma}(\mathbf{D}_{F}),\psi_{\sigma}(\mathbf{D}_{M}(\phi_{\theta})))\!+\!\lambda\mathcal{L}_{reg}(\phi_{\theta}),\end{array}$ (5)
where $\phi^{*}$ denotes the optimal displacement field, the $\psi_{\sigma}$ is a Gaussian smoothing function with variance $\sigma$, $\mathcal{L}_{sim}(\mathbf{D}_{F},\mathbf{D}_{M}(\phi_{\theta}))$ denotes the negated cosine similarity function that quantifies the dissimilarity between the $\mathbf{D}_{F}$ and the $\mathbf{D}_{M}(\phi_{\theta})$ spatially transformed by $\phi$, and $\mathcal{L}_{reg}(\phi)=||\nabla\phi||^{2}_{2}$ represents the smoothness regularization function that penalizes implausible solution. Our proposed DNS is differentiable and can be adapted to iterative instance optimization methods [40, 51] and learning-based registration frameworks [6, 37, 35, 36]. To exemplify the flexibility of DSIR, we parameterize the registration function with one learning-based method, LapIRN [37], and one 3-level non-rigid instance optimization method (IO) [38], with implementation details shown in the supplementary material.
@@ 4.1 Data and Pre-processing
We collected a 3D multi-phase contrast-enhanced liver CT dataset from hospitals in China to evaluate the inter-patient registration. This dataset consists of 1966 patients with tumors, and each patient contains scans in three phases, i.e., pre-contrast, arterial and venous phases, acquired from the CT scanner at different time points after the contrast agent injection. There are large non-linear deformations and heterogeneous image intensities across multi-phase CT due to gravity, muscle contractions, and contrast agent injection. We divided the dataset into 1886/10/70 cases for training, validation, and test sets. The liver and tumour masks of the validation and test sets are manually annotated and checked by a senior radiologist.
@@
We evaluated our method on the Abdomen MR-CT task of the Learn2Reg challenge 2021 [22]. The dataset comprises 8 sets of paired MR and CT scans and 90 unpaired CT/MR scans, both depicting the abdominal region of a single patient and presenting notable deformations. Since there are only limited paired MR and CT scans, we only evaluate the iterative optimization-based registration methods in this task. We use all 8 paired MR and CT scans as the test set. Each scan has four manually labeled anatomical structures: liver, spleen, right kidney, and left kidney. For MASR-Net, we split the 90 unpaired CT/MR scans into 85 and 5 scans for training and validation sets.
@@
We performed atlas-based registration on BraTS18 https://www.med.upenn.edu/sbia/brats2018.html and ISeg19 https://iseg2019.web.unc.edu/ datasets [34, 53]. The training set consists of 135 cases, each with T1-weighted contrast (T1w) and T2-weighted contrast (T2w) images, and was resampled to 1mm3 isotropic resolution. 10 cases contain full delineation of anatomical structures, including grey matter, white matter and cerebrospinal fluid (CSF) for both T1w and T2w modalities. We randomly select 1 case from the test set as the atlas and register the remaining 9 cases to the atlas. Similar to Abdomen MR-CT, we only evaluate iterative optimization-based registration methods on this task due to the limited training data.
@@ 4.3 Evaluation metrics and Baselines
We compare our method with three state-of-the-art conventional registration methods (ANTs [4], NiftyReg [52] and DEEDs [18]) and two learning-based approaches (Voxelmorph [6] and LapIRN [41]). Specifically, we use the ANTs registration implementation in the publicly available ANTs software package [3]. Both ANTs and NiftyReg methods use a four-level multi-resolution strategy with adaptive gradient descent optimization and MI as the similarity measure. For Voxelmorph and LapIRN, we use official implementations and the best hyperparameter reported in their paper. By default, all learning-based methods are trained with two similarity metrics (NMI and MIND).
@@ 4.4 Results
Flexibility and versatility. Compared to the results of LapIRN with different metrics, LapIRN with our DNS achieves the best overall registration performance in tumor and organ alignment, showing +2.22 and +0.75 DSC gains over the one with MIND in tumor and organ alignment, respectively. More importantly, when comparing the results of IO with DNS and MIND, there are significant improvements in registration accuracy of both tumor and organ (+4.16 and + 1.72 gains in DSC). The result proves that our method is more robust and flexible with diverse registration frameworks than MIND. When using DNS with a simple iterative gradient optimization, the DSIR of DNS is more discriminative than MIND and can lead to less ambiguous matching during the iterative optimization. Besides, although NMI shows comparable DSC in multiphase CT registration, the qualitative results in Fig. 3 demonstrate that methods using MI/NMI fail to recover the large non-linear misalignment in MR-CT registration.
@@
Robust to large deformation. The initial DSC of this task is relatively low (37.32), indicating there are possible large spatial misalignments between image pairs. All conventional methods (ANTs and NiftyReg) with MI fail spectacularly in this task, suggesting that MI cannot register multimodal images with large deformation. In contrast to MI, structural image representation-based methods (DEEDs and DNS) achieve decent results in this task (83.53 and 85.61 DSC, respectively). It is worth noting that our DNS with a solely simple gradient descent-based optimization strategy (IO) outperforms DEEDs, which employ a more sophisticated convex optimization.
@@ 4.4.1 Ablation Studies
Effect of Deep Neighbourhood Self-similarity and contrastive learning. Tab. 3 demonstrates performance metrics for different registration models on the multiphase CT registration task (Pre-contrast $\leftarrow$ Venous & Arterial). Although a recent study [13] reported that random features from CNN could be expressive, the DSIR of random initialize CNN has lower DSC than that of initial, suggesting that random initialization alone is insufficient for multi-modal registration. Moreover, there is a significant performance gain when the DNS module is added to the randomly initialized CNN (+11.12 DSC), suggesting the effectiveness of DNS in extracting deep structural information. Comparing the feature similarity analysis of random initialization (MASR-Net without training) and DNS in Fig. 4, one can observe that the proposed anatomy-aware contrastive learning effectively improves the discriminability of the DSIR of DNS, resulting in high feature similarity centralized in the true anatomical correspondence and relatively low feature similarity at the other anatomical structures.
@@ 5 Conclusion
This paper proposes a deep structural image representation learning method dedicated to multi-modality medical image registration. Our method leverages deep neighbourhood self-similarity to learn highly discriminative, contrast invariance structural representations for multimodal images. Anatomy-aware contrastive learning is introduced to further enhance the expressiveness and discriminability of the learned structural representation, reducing the ambiguity in matching the true anatomical correspondence between multimodal images. Comprehensive experiments demonstrate that our method achieves state-of-the-art registration accuracy compared with conventional local structural representations without the need for a dedicated optimizer.